Circular No. 6266

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 B. A. Harmon, W. S. Paciesas, S. N. Zhang and K. J. Deal, NASA/

Marshall Space Flight Center, report, for the Compton Observatory BATSE Team: “Hard x-ray flux (20–100 keV) from GRS 1915+105 has increased over the past month to a level comparable to the intense outbursts seen in 1992 and 1994 (IAUC 5590, 5959). The flux reached 300 mCrab around Oct. 15 and has remained near this level (within 50 percent) as of Nov. 20. Less intense outbursts reaching about 200 mCrab have also been observed since June 1995. These are nonperiodic but are of similar intensity and each about 15–20 days in duration. The source continues to be variable on a daily basis. The hard x-ray activity correlates well on long timescales (days to weeks) with the radio flux, but no strong signature is seen in xrays that corresponds to the shorter-timescale (hours to days) radio flares reported below.”

E. B. Waltman and R. S. Foster, Naval Research Laboratory; F. D. Ghigo, National Radio Astronomy Observatory; and M. Tavani, Columbia University, report: “The source GRS 1915+105 has been active at 2.25 and 8.3 GHz during the time period Oct. 10–Nov. 20 as monitored by the Green Bank Interferometer. Radio flares as large as 200–400 mJy (2.25 GHz) have been observed four times. Flux densities between these flares ranged from 10 to 100 mJy, compared to a quiescent level of 10 mJy or less prior to the activity that began on Aug. 10 (IAUC 6204).”

4U 0115+634 M. H. Finger, M. Scott and K. Hagedon, Universities Space Research

Association; R. B. Wilson and C. A. Wilson, NASA/Marshall Space Flight Center; M. Stollberg and W. S. Paciesas, University of Alabama, Huntsville; and T. A. Prince and B. Vaughan, California Institute of Technology, report for the BATSE/Compton GRO Observatory pulsar team: “Pulsed hard xray flux is currently being detected from the transient x-ray pulsar 4U 0115+634. The source was detected on Nov. 18, when the rms pulsed flux (20–50 keV) was ∼ 15 mCrab. This is the first detection by BATSE of this source since the outburst that occurred during 1994 May 8–June 27 (IAUC 5990, 5999). On 1995 Nov. 21 the rms pulsed flux (20–50 keV) was ∼ 70 mCrab. The pulse frequencies observed from Nov. 19 to 21 are consistent with a neutron-star spin rate of 0.2766635 ± 0.0000003 Hz at epoch JD 2450042 and a spin-up rate of 1.4 ± 0.3 × 10^{−11} Hz/s, using the binary orbital ephemeris of Cominsky et al. (1994, 2nd Compton Symposium, AIP Proc. 304, 294).”

1995 November 22 Brian G. Marsden